Mathematics 134 – Intensive Calculus for Science 2 Quiz 7 – March 31, 2006

Your Name:

Do all work on this sheet (front and back if needed). You have 15 minutes. No calculators on this quiz.

A) (10) Find the sum of the infinite geometric series:

7 − \frac{7}{3} + \frac{7}{9} − \frac{7}{27} + ⋯

B) (15) Find the Taylor polynomial of degree n = 3 for the function f(x) = (1 + 2x)^{1/3} at a = 0.